# Computation method for perfectoid purity and perfectoid BCM-regularity

### v5

@@ 1. Introduction
$\inf\left\{i/p^{e}\geq 0\middle|\begin{array}[]{l}\text{there exist a perfectoid $R$-algebra $B$ a $p^{e}$-th root $\varpi$}\\ \text{of $p$, such that the map $R\to B\xrightarrow{\cdot\varpi^{i}}B$ is pure}\end{array}\right\},$
@@
We further assume that $\operatorname{Spec}{\overline{S}}\backslash\{\overline{S}_{>0}\}$ is strongly $F$-regular. If $S_{(p,S_{>0})}$ is quasi-$F$-split, then $S_{(p,S_{>0})}$ is perfectoid BCM-regular. Furthermore, the converse implication holds if $p=2$.
@@ 2.1. Notation
    For a ring homomorphism $f\colon A\to B$ and a finitely generated ideal $I$ of $A$, we say that $f$ is $I$-completely faithfully flat if for every $I$-torsion $A$-module $N$, we have $\operatorname{Tor}_{i}^{A}(B,N)=0$ for every $i\neq 0$ and $B/IB$ is faithfully flat over $A/I$. If $I=(p)$, we say that $f$ is $p$-completely faithfully flat.
@@ 2.2. The ring of Witt vectors
    Then $W(A)$ is a ring with $1=(1,0,\ldots)\in W(A)$, which is called the ring of Witt vectors over $A$. For $a\in A$, we set $[a]:=(a,0,\ldots)\in W(A)$. We define an additive map $V$ by
@@
    We note that $V^{n}W(A)\subseteq W(A)$ is an ideal. For $a\in A$ and $n\in\mathbb{Z}_{\geq 1}$, we set $[a]:=(a,0,\ldots,0)\in W_{n}(A)$. We define a $W(A)$-module homomorphism $R$ by
@@ 2.3. Test perfectoid and perfectoid pure threshold
where ()(*)( ∗ ) follows from the faithfully flatness of $R_{\mathrm{perf}}\to B$. ∎
@@
$\inf\left\{i/p^{e}\geq 0\middle|\begin{array}[]{l}\textup{there exist a perfectoid $R$-algebra $B$ and $\varpi$ a $p^{e}$-power}\\ \textup{root of $p$ such that $R\to B\xrightarrow{\cdot\varpi^{i}}B$ is pure}\end{array}\right\}.$
@@
[…]
@@ 3. Witt vectors of rings with Frobenius lifts
=$\overset{(\star)}{=}V(s_{\phi}\circ\phi(b)\Delta_{W_{n}}(\alpha)+s_{\phi}\circ\phi(a)\Delta_{W_{n}}(\beta)+p\Delta_{W_{n}}(\alpha)\Delta_{W_{n}}(\beta)),$
where ()($\star$)( ⋆ ) follows from Proposition 3.3 and Proposition 2.3 (2). Therefore, we have
@@
$[a]=s_{\phi}(a)+Vs_{\phi}(\Delta_{1}(a))+\cdots+V^{n-1}s_{\phi}(\Delta_{n-1}(a))+V^{n}\Delta_{n}(a).$
@@
$[a]=s_{\phi}(a)+Vs_{\phi}(\Delta_{1}(a))+\cdots+V^{n-1}s_{\phi}(\Delta_{n-1}(a))+V^{n}\Delta_{n}(a).$
@@
=$\overset{(\star_{1})}{=}\sum^{n}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)-\sum^{n-1}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)$
@@ 4.1. Definition of quasi-$F$-splitting
is well-defined and a ring homomorphism, where $\tilde{a}\in A$ is a lift of $a\in A$. We take $a,b\in A$ such that $\pi(a)=0$. Then we have $[(a+pb)^{p}]\equiv[a^{p}]\ \textrm{mod}\ p$ and $[a^{p}]$ is contained in the kernel of $W_{n}(\pi)$. Thus, $\Phi_{A,n}$ induces the ring homomorphism $\Phi_{R,n}$, as desired. ∎
@@
First, we prove that $V^{m}\colon F^{m}_{*}Q_{R,n}\to Q_{R,n+m}$ is an $\overline{R}$-module homomorphism. We take $\alpha\in W_{n}(R)$ and $b\in R$. Then we have $[b^{p}]V^{m}\alpha=V([b^{p^{m+1}}]\alpha)$ by Proposition 2.3, thus $V^{m}$ is an $\overline{R}$-module homomorphism. Next, we prove the exactness. It is enough to show that the sequence
@@
[…]
@@
We use the notation introduced in Notation 4.2 and assume $(R,\mathfrak{m})$ is Noether local $p$-torsion free and $\overline{R}$ is $F$-finite. Then we have $\operatorname{ht}(R)=\operatorname{ht}(\widehat{R})$, where $\widehat{R}$ is the $m$-adic completion of $R$. Furthermore, $R$ is quasi-$(F,F^{\infty})$-split if and only if so is $\widehat{R}$.
@@
[…]
@@
[…]
where $d:=\dim\overline{R}$ and the injectivity of ()($\star$)( ⋆ ) follows from Proposition 4.8. Since the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ is contained in the image of the map $H^{d-1}_{\mathfrak{m}}(\overline{R/f})\to H^{d}_{\mathfrak{m}}(\overline{R})$, we obtain that $\Phi_{R/f,n}$ is injective if and only if so is $\Phi^{f}_{R,n}$, as desired. ∎
@@ 4.2. Quasi-$F$-splitting for perfectoids
In this subsection, we study the purity of $\Phi_{R,n}$ for a perfectoid $R$. Theorem 4.10 is related to the computation of perfectoid pure threshold (cf. LABEL:ht-to-p-pure).
@@
Let $R$ be a $p$-torsion free perfectoid and $n$ a positive integer. We assume $R$ has a compatible system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$. Then there exist an $\overline{R}$-module homomorphism $\overline{\psi}_{n}\colon Q_{R,n}\to R/(p^{i_{n}})$ and a unit $w\in\overline{R}$ such that $\overline{\psi}_{n}(1)=1$ and $\overline{\psi}_{n}(V^{n-1}[a])=wp^{i_{n-1}}F^{-n}(a)$ for every $a\in R$, where
@@
Let $R^{\flat}$ be a tilt of $R$ and $T\in R^{\flat}$ the element corresponding to a system $\{p^{1/p^{e}}\}$. Let $A:=W(R^{\flat})$ and $d:=[T]-p$, then $(A,(d))$ is a perfect prism with $A/(d)\simeq R$ by [BS]*Corollary 2.31. The lift of Frobenius on $A$ is denoted by $\phi$. We prove the following claim:
@@
=$\overset{(\star_{1})}{=}$ $T^{1/p}\psi^{\prime}(\Delta_{W_{n}}([a_{0}]))+\psi^{\prime}((a_{1}d,\ldots,a_{n-1}d))\overset{(\star_{2})}{\in}(T^{i_{n}}),$
@@ 5.1. Quasi-$F$-splitting versus perfectoid purity
[…]
@@
the element $\eta$ maps to 00. Since $R$ is $p$-torsion free and Gorenstein, the first homomorphism in (5.1) is injective, thus we obtain that the map
@@
We use the notation introduced in Notation 5.1 and assume that $R$ is quasi-$F$-split. We take $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ and construct sequences $\{\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}\}_{i\geq 0}$ and $\{n_{i}\in\mathbb{Z}_{\geq 1}\}_{i\geq 0}$ using the following procedure. We start with $\eta_{0}:=\eta$. Suppose that $\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ has been defined for some $i\geq 0$. We take the minimal integer $n_{i}\geq 1$ such that $\Phi_{R,n_{i}}(\eta_{i})\neq 0$. We note that such an integer $n_{i}$ exists since $R$ is quasi-$F$-split and $n_{i}\leq\operatorname{ht}(R)$. By the exact sequence in Proposition 4.3, there exists $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{R})$ such that $V^{n_{i}-1}(\eta^{\prime})=\Phi_{R,n_{i}}(\eta_{i})$. We choose such an $\eta^{\prime}$ and set $\eta_{i+1}:=\eta^{\prime}$.
@@
Let $(A,\mathfrak{m})$ be a local Gorenstein $\mathbb{Z}_{(p)}$-algebra such that $A$ is $p$-torsion free and $p\in\mathfrak{m}$. Let $f\in\mathfrak{m}$ be an element of $A$ such that $f,p$ is a regular sequence of $A$. Let $A_{\infty}$ be a perfectoid over $A$ such that $(A_{\infty}/(p))_{perfd}\simeq(A/(p))_{perf}$, $A_{\infty}$ is a cohomologically Cohen-Macaulay, $A\to A_{\infty}$ is pure, $A_{\infty}$ has a system of $p$-power roots system of $f$, and $f,p$ are regular elements of $A_{\infty}$. Let $R:=A/(f)$ and $R_{\infty}:=(A_{\infty}/(f))_{perfd}$. If $R$ is quasi-$F$-split, then $R\to R_{\infty}$ is pure.
@@
$[a^{p}]\equiv V^{n-1}(b)\ \textrm{mod}\ ([x_{1}],\ldots,[x_{d}])Q_{R,n}.$
@@
$[a^{p}]+(c_{0}f,\ldots,c_{n-1}f)\equiv V^{n-1}(b)\ \textrm{mod}\ ([x_{1}],\ldots,[x_{d}])Q_{A,n}.$
@@
[…]
@@ 5.2. Computation of perfectoid threshold via quasi-$F$-splitting
[…]
@@
[…]
@@
(cf. [p-pure]*Proposition 6.7) Let $(R,\mathfrak{m})$ be a Noetherian complete local domain with $p\in\mathfrak{m}\backslash\{0\}$. We assume $R$ is complete intersection. If $R/pR$ is quasi-$F$-split and $R[1/p]$ is regular, then $R$ is BCM-regular.

### v7

@@ 1. Introduction
[…]
@@
[…]
@@ 2.1. Notation
    For a ring homomorphism $f\colon A\to B$ and a finitely generated ideal $I$ of $A$, we say that $f$ is $I$-completely faithfully flat if for every $I$-torsion AA-module NN, we have $\operatorname{Tor}_{i}^{A}(B,N)=0$ for every $i\neq 0$ and $B/IB$ is faithfully flat over $A/I$. If $I=(p)$, we say that $f$ is $p$-completely faithfully flat.
@@ 2.2. The ring of Witt vectors
    Then $W(A)$ is a ring with $1=(1,0,\ldots)\in W(A)$, which is called the ring of Witt vectors over $A$. For $a\in A$, we set $a]:=(a,0,\ldots)\in W(A)$. We define an additive map $V$ by
@@
    We note that $V^{n}W(A)\subseteq W(A)$ is an ideal. For $a\in A$ and $n\in\mathbb{Z}_{\geq 1}$, we set $a]:=(a,0,\ldots,0)\in W_{n}(A)$. We define a $W(A)$-module homomorphism $R$ by
@@ 2.3. Test perfectoid and perfectoid pure threshold
where ()(*) follows from the faithfully flatness of $R_{\mathrm{perf}}\to B$. ∎
@@
[…]
@@
${R}$${R_{\infty}}$${R_{\infty}}$${B}$${B,}$$\scriptstyle{\cdot p^{1/p^{e}}}$$\scriptstyle{\cdot\nu(p^{1/p^{e}})}$
@@ 3. Witt vectors of rings with Frobenius lifts
$\overset{(\star)}{=}V(s_{\phi}\circ\phi(b)\Delta_{W_{n}}(\alpha)+s_{\phi}\circ\phi(a)\Delta_{W_{n}}(\beta)+p\Delta_{W_{n}}(\alpha)\Delta_{W_{n}}(\beta)),$
where $(\star)$ follows from Proposition 3.3 and Proposition 2.3 (2). Therefore, we have
@@
$a]=s_{\phi}(a)+Vs_{\phi}(\Delta_{1}(a))+\cdots+V^{n-1}s_{\phi}(\Delta_{n-1}(a))+V^{n}\Delta_{n}(a).$
@@
$a]=s_{\phi}(a)+Vs_{\phi}(\Delta_{1}(a))+\cdots+V^{n-1}s_{\phi}(\Delta_{n-1}(a))+V^{n}\Delta_{n}(a).$
@@
$\overset{(\star_{1})}{=}\sum^{n}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)-\sum^{n-1}_{s=0}p^{s}\phi^{n-s}\Delta_{s}(a)$
@@ 4.1. Definition of quasi-$F$-splitting
is well-defined and a ring homomorphism, where $\tilde{a}\in A$ is a lift of $a\in A$. We take $a,b\in A$ such that $\pi(a)=0$. Then we have $(a+pb)^{p}]\equiv[a^{p}]\ \textrm{mod}\ p$ and $a^{p}]$ is contained in the kernel of $W_{n}(\pi)$. Thus, $\Phi_{A,n}$ induces the ring homomorphism $\Phi_{R,n}$, as desired. ∎
@@
First, we prove that $V^{m}\colon F^{m}_{*}Q_{R,n}\to Q_{R,n+m}$ is an $\overline{R}$-module homomorphism. We take $\alpha\in W_{n}(R)$ and $b\in R$. Then we have $b^{p}]V^{m}\alpha=V([b^{p^{m+1}}]\alpha)$ by Proposition 2.3, thus $V^{m}$ is an $\overline{R}$-module homomorphism. Next, we prove the exactness. It is enough to show that the sequence
@@
${F^{e}_{*}\overline{R}}$${F^{n+e-1}_{*}\overline{R}}$${Q_{R,n}}$${Q^{e}_{\overline{R},n},}$$\scriptstyle{F^{e}}$$\scriptstyle{V^{n-1}}$$\scriptstyle{V^{n-1}}$$\scriptstyle{F^{e}}$
@@
We use the notation introduced in Notation 4.2 and assume $(R,\mathfrak{m})$ is Noether local $p$-torsion free and $\overline{R}$ is $F$-finite. Then we have $ht(R^)\operatorname$$\operatorname{ht}(\widehat{R})$, where $\widehat{R}$ is the $m$-adic completion of $R$. Furthermore, $R$ is quasi-$(F,F^{\infty})$-split if and only if so is $\widehat{R}$.
@@
${0}$${\overline{R}}$${\overline{R}}$${\overline{R/f}}$${0}$${0}$${Q^{f}_{R,n}}$${Q_{R,n}}$${Q_{R/f,n}}$${0,}$$\scriptstyle{\cdot f}$$\scriptstyle{\Phi^{f}_{R,n}}$$\scriptstyle{\Phi_{R,n}}$$\scriptstyle{\Phi_{R/f,n}}$
@@
${0}$${H^{d-1}_{\mathfrak{m}}(\overline{R/f})}$${H^{d}_{\mathfrak{m}}(\overline{R})}$${H^{d}_{\mathfrak{m}}(\overline{R})}$${0}$${0}$${H^{d-1}_{\mathfrak{m}}(Q_{R/f,n})}$${H^{d}_{\mathfrak{m}}(Q^{f}_{R,n})}$${H^{d}_{\mathfrak{m}}(Q_{R,n})}$${0,}$$\scriptstyle{\Phi_{R/f,n}}$$\scriptstyle{\cdot f}$$\scriptstyle{\Phi^{f}_{R,n}}$$\scriptstyle{\Phi_{R,n}}$$\scriptstyle{(\star)}$
where $d:=\dim\overline{R}$ and the injectivity of $(\star)$ follows from Proposition 4.8. Since the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ is contained in the image of the map $H^{d-1}_{\mathfrak{m}}(\overline{R/f})\to H^{d}_{\mathfrak{m}}(\overline{R})$, we obtain that $\Phi_{R/f,n}$ is injective if and only if so is $\Phi^{f}_{R,n}$, as desired. ∎
@@ 4.2. Quasi-$F$-splitting for perfectoids
In this subsection, we study the purity of $\Phi_{R,n}$ for a perfectoid $R$. Theorem 4.10 is related to the computation of perfectoid pure threshold (cf. Theorem 5.11).
@@
(cf. [BMS18]*Lemma 3.12 Let $R$ be a $p$-torsion free perfectoid and $n$ a positive integer. We assume $R$ has a compatible system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$. Then there exist an $\overline{R}$-module homomorphism $\overline{\psi}_{n}\colon Q_{R,n}\to R/(p^{i_{n}})$ and a unit $w\in\overline{R}$ such that $\overline{\psi}_{n}(1)=1$ and $\overline{\psi}_{n}(V^{n-1}[a])=wp^{i_{n-1}}F^{-n}(a)$ for every $a\in R$, where
@@
Let $R^{\flat}$ be a tilt of $R$ and $T\in R^{\flat}$ the element corresponding to a system $\{p^{1/p^{e}}\}$. Let $A:=W(R^{\flat})$ and d:=[T]pd:=[T]-p, then $(A,(d))$ is a perfect prism with $A/(d)\simeq R$ by [BS]*Corollary 2.31. The lift of Frobenius on $A$ is denoted by $\phi$. We prove the following claim:
@@
$\overset{(\star_{1})}{=}$ $T^{1/p}\psi^{\prime}(\Delta_{W_{n}}([a_{0}]))+\psi^{\prime}((a_{1}d,\ldots,a_{n-1}d))\overset{(\star_{2})}{\in}(T^{i_{n}}),$
@@
Remark 4.12.
By [BMS18]*Lemma 3.12, the homomorphism $\theta_{n}$ induces the isomorphism
$\overline{\theta_{n}}\colon R^{\flat}/((p^{\flat})^{1+1/p+\cdots+1/p^{n-1}})\xrightarrow{\sim}\overline{W}_{n}(R).$
Consider the composition
$Q_{R,n}\xrightarrow{\overline{\theta_{n}}^{-1}}F_{*}R^{\flat}/((p^{\flat})^{1+1/p+\cdots+1/p^{n-1}})\xrightarrow{F^{-1}}R^{\flat}/((p^{\flat})^{(1/p+\cdots+1/p^{n})})\simeq R/(p^{i_{n}}),$
then we get a splitting of $\Phi_{R,n}$.
@@ 5.1. Quasi-$F$-splitting versus perfectoid purity
${0}$${R_{\infty}}$${R_{\infty}}$${\overline{R}_{\infty}}$${0}$${0}$${R_{\infty}}$${R_{\infty}}$${\overline{R}_{\infty}/(p^{i_{n}})}$${0,}$$\scriptstyle{\cdot p}$$\scriptstyle{\cdot p^{1-i_{n}}}$$\scriptstyle{\cdot p^{i_{n}}}$
@@
the element $\eta$ maps to 0. Since $R$ is $p$-torsion free and Gorenstein, the first homomorphism in (5.1) is injective, thus we obtain that the map
@@
We use the notation introduced in Notation 5.1 and assume that $R$ is quasi-$F$-split. We take $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ and construct sequences $\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}\}_{i\geq 0}$ and $\{n_{i}\in\mathbb{Z}_{\geq 1}\}_{i\geq 0}$ using the following procedure. We start with $\eta_{0}:=\eta$. Suppose that $\eta_{i}\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ has been defined for some $i\geq 0$. We take the minimal integer $n_{i}\geq 1$ such that $\Phi_{R,n_{i}}(\eta_{i})\neq 0$. We note that such an integer $n_{i}$ exists since $R$ is quasi-$F$-split and $n_{i}\leq\operatorname{ht}(R)$. By the exact sequence in Proposition 4.3, there exists $\eta^{\prime}\in H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{R})$ such that $V^{n_{i}-1}(\eta^{\prime})=\Phi_{R,n_{i}}(\eta_{i})$. We choose such an $\eta^{\prime}$ and set $\eta_{i+1}:=\eta^{\prime}$.
@@
Let $(A,\mathfrak{m})$ be a local Gorenstein $\mathbb{Z}_{(p)}$-algebra such that $A$ is $p$-torsion free and $p\in\mathfrak{m}$. Let $f\in\mathfrak{m}$ be an element of $A$ such that $f,p$ is a regular sequence of $A$. Let $A_{\infty}$ be a perfectoid over $A$ such that $A_{\infty}$ is a cohomologically Cohen-Macaulay, $A\to A_{\infty}$ is pure, $A_{\infty}$ has a system of $p$-power roots system of $f$, and $f,p$ are regular elements of $A_{\infty}$. Let $R:=A/(f)$ and $R_{\infty}:=(A_{\infty}/(f))_{perfd}$. If $R$ is quasi-$F$-split, then $R\to R_{\infty}$ is pure.
@@
$a^{p}]\equiv V^{n-1}(b)\ \textrm{mod}\ ([x_{1}],\ldots,[x_{d}])Q_{R,n}.$
@@
$a^{p}]+(c_{0}f,\ldots,c_{n-1}f)\equiv V^{n-1}(b)\ \textrm{mod}\ ([x_{1}],\ldots,[x_{d}])Q_{A,n}.$
@@
${0}$${H^{d+1}_{\mathfrak{m}}(\overline{A})}$${H^{d+2}_{\mathfrak{m}}(A)}$${0}$${H^{d+1}_{\mathfrak{m}}(\overline{A}_{\infty})}$${H^{d+2}_{\mathfrak{m}}(A_{\infty}).}$$\scriptstyle{\alpha}$$\scriptstyle{\cdot f^{1-1/p^{e}}}$$\scriptstyle{\cdot f^{1-1/p^{e}}}$
@@ 5.2. Computation of perfectoid threshold via quasi-$F$-splitting
${\eta_{j}}$${H^{d}_{\mathfrak{m}}(\overline{R})}$${H^{d}_{\mathfrak{m}}(Q_{R,n})}$${H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{R})}$${\eta_{j+1}}$${\eta^{\prime}_{j}}$${H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})}$${H^{d}_{\mathfrak{m}}(Q_{R_{\infty},n})}$${H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{R}_{\infty})}$${\eta^{\prime}_{j+1}}$${H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{i_{n}}))}$${H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{1/p^{n}})),}$${F^{-n}(\eta^{\prime}_{j+1})}{\in}$$\scriptstyle{\Phi_{R,n}}$$\scriptstyle{V^{n-1}}{\ni}{\in}$$\scriptstyle{\Phi_{R_{\infty},n}}$$\scriptstyle{\alpha}$$\scriptstyle{\overline{\psi}_{n}}$$\scriptstyle{F^{-n}}$$\scriptstyle{V^{n-1}}{\ni}$$\scriptstyle{\cdot wp^{i_{n}-1}}{\ni}$
@@
${H^{d+1}_{\mathfrak{m}}(R_{\infty})}$${H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})}$${H^{d}_{\mathfrak{m}^{\flat}}(R_{\infty}^{\flat}/(T))}$${H^{d+1}_{\mathfrak{m}^{\flat}}(R^{\flat}_{\infty}),}\scriptstyle{\sim}$
@@
[…]